An elevator weighing 12 kN starts from rest and acquires an upward velocity of 2 m/sec in a distance of 5 m. If the acceleration is constant, what is the tension in the cable? *


1
Expert's answer
2022-02-14T14:36:30-0500

We know that

Newton motion third

Law

"V^2=U^2+2aS"

"a=\frac{V^2-U^2}{2S}"

Put value

"a=\frac{2^2-0^2}{2\times5}=0.4m/sec^2"

Now We know that

"T-ma=mg"

"T=mg(1+\frac{a}{g})=12000\times(1+\frac{0.4}{9.8})=12.489KN"